The sum of two consecutive even integers, of which the smaller is x, is no more than 24.
Answer by stanbon(75887)

The sum of two consecutive even integers,
of which the smaller is x, is no more than 24.
-----
Note: Even integers are not represented by x.
Even integers are ALWAYS a multiple of 2, so they
must be represented by 2*x or 2(x+1) etc.
--------
1st: 2x
2nd: 2x+2
--------
Equation::
2x + 2x+2 <= 24
x + x+1 <= 12
2x <= 11
x <= 5.5
----------------
2x <= 11
2x+2 <= 13
-----------------------
Cheers,
Stan H.
--------------------